# Pushing the Envelope of Low-Bit LLM via Dynamic Error Compensation

### v1

@@ Abstract
Quantization of Large Language Models (LLMs) has recently gained popularity, particularly for on-device settings with limited hardware resources. While efficient, quantization inevitably degrades model quality, especially in aggressive low-bit settings such as 3-bit and 4-bit precision. In this paper, we propose QDEC, an inference scheme that improves the quality of low-bit LLMs while preserving the key benefits of quantization: GPU memory savings and inference latency reduction. QDEC stores the residual matrix—the difference between full-precision and quantized weights—in CPU, and dynamically fetches the residuals for only a small portion of the weights. This portion corresponds to the salient channels, marked by activation outliers, with the fetched residuals helping to correct quantization errors in these channels. Salient channels are identified dynamically at each decoding step by analyzing the input activations—this allows for the adaptation to the dynamic nature of activation distribution, and thus maximizes the effectiveness of error compensation. We demonstrate the effectiveness of QDEC by augmenting state-of-the-art quantization methods. For example, QDEC reduces the perplexity of a 3-bit Llama-3-8B-Instruct model from 10.15 to 9.12—outperforming its 3.5-bit counterpart—while adding less than 0.0003% to GPU memory usage and incurring only a 1.7% inference slowdown on NVIDIA RTX 4050 Mobile GPU. The code will be publicly available soon.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) based on the Transformer architecture [58] have shown great potential to reshape various aspects of our daily lives [1, 55, 43, 39, 54]. However, their deployment costs, largely due to their substantial size, present a significant barrier to broader accessibility. The large model size increases memory requirements, driving up system costs, and also results in longer inference latencies, thereby limiting potential use cases [59].
Quantization has gained attention as a promising solution for mitigating the deployment costs of LLMs [68, 27]. By lowering the precision of model parameters, quantization addresses both memory limitations and inference latency. The importance of quantization is especially pronounced for on-device deployments, where model compression is often a necessity rather than an option due to strict memory budgets. In such scenarios, it is a common practice to explore the trade-off between model size and quality by adjusting the quantization levels, in order to find the most effective solution within the given memory constraints [15, 14, 60, 8, 57].
While quantization may allow gigantic LLMs to fit into small-memory devices—something that would have been impossible otherwise—it often leads to model quality degradation due to the inevitable loss of information. This is especially true for low-bit settings, such as 3-bit and 4-bit quantization, which are often used to accommodate the parameter sizes of LLMs [30, 35, 19]. This raises a key research question that this paper addresses: given a quantized LLM configured with the best possible effort under the memory budget, is there a way to recover the quality loss caused by quantization?
@@
A clue for this problem comes from the well-known fact that not all channels in the weight matrix are equally important for weight quantization. Some channels, referred to as salient channels, are more critical, primarily due to the presence of activation outliers [10, 7, 24]. When certain input activation values are particularly large, the quantization errors in the corresponding weight channels—those multiplied by these large activation values—are amplified, making the channels salient. By identifying these channels and selectively fetching error compensation terms from CPU memory, we can maximize the quality boost while minimizing data transfer.
Hence, the accurate identification of these salient channels is essential for overcoming the limited communication bandwidth, allowing for the transfer of only impactful information. Previous works that attempt to improve quantization algorithms by addressing activation outliers often analyze the distribution of activation values on a calibration dataset to predetermine the salient channels [10, 32, 26, 35]. However, this approach is suboptimal because it statically designates certain weight channels as salient throughout the inference, while the real distribution of activation values—and thus the distribution of salient channels—changes dynamically at each decoding step. Accounting for the dynamic nature of activation outliers is crucial for the accurate identification of salient channels.
Thus, in this paper, we introduce QDEC (Quantization with Dynamic Error Compensation), an inference scheme for quantized LLMs that dynamically identifies salient channels and compensates for quantization errors in these channels, in real time. QDEC enhances model quality while fully preserving the two main benefits of quantization, namely reduced GPU memory usage and improved inference latency, by efficiently utilizing CPU memory. To achieve this, QDEC stores the residuals of the quantized weight matrices in CPU memory, fetching only the parts that correspond to the dynamically identified salient channels for error compensation. Although small in size, these residuals provide a significant quality boost. This dynamic error compensation is performed concurrently with inference by a highly optimized GPU kernel, ensuring that all additional operations are seamlessly integrated into the existing workflow, thereby minimizing any inference slowdown. Below are the key contributions of our work:
@@
  • We present QDEC, an inference scheme that enhances quantized LLMs by dynamically identifying salient channels and compensating quantization errors in them.
  • We introduce a parameter tuner for QDEC that automatically recommends system parameters to satisfy a target latency bound.
  • We provide a comprehensive evaluation of QDEC across five different consumer-grade GPUs, demonstrating significant quality improvement with minimal memory and inference latency overhead.
@@ 2.2 LLM Quantization
Quantization for LLMs—a popular compression technique that reduces both memory usage and inference latency—can be categorized into two main types [68, 27]: weight-activation quantization [64, 10, 48, 63, 36, 33] and weight-only quantization [19, 35, 30, 12, 32, 9, 56, 17, 11, 29]. Each are suited to different inference scenarios. Weight-activation quantization is primarily used in datacenter settings, where both memory and computational costs must be minimized to improve throughput. Quantizing both weights and activations allows for the efficient use of low-precision arithmetic units (e.g., INT4, INT8, FP8) available on modern GPUs [2]. In contrast, for on-device inference, weight-only quantization is the preferred approach [68, 30]. In this apporach, quantized weights are loaded from memory and dequantized on-the-fly to full precision (i.e., FP16), before being multiplied with the full-precision activations [44]. While this reduces only the memory bandwidth and not the computation, the memory reduction alone is sufficient to achieve speedups in on-device inference, where memory bandwidth is the primary bottleneck, as discussed in Section 2.1.
Weight-only quantization can be further divided into two sub-categories: quantization-aware training (QAT) [36, 11, 29], and post-training quantization (PTQ) [19, 35, 30, 12, 32, 9, 56, 17]. While QAT, which mitigates quantization errors through retraining, typically yields better results, the high costs of retraining make it impractical for many end-users [30, 68]. As a result, PTQ, which does not require retraining, has become the preferred quantization method for on-device LLM inference. Therefore, in this paper, we specifically focus on weight-only PTQ for on-device inference.
@@ 3 Augmenting Quantized LLM with CPU Memory
In this section, we propose leveraging CPU memory—an often underutilized resource in LLM inference, as GPUs are the de facto standard processors for this workload—as a means to augment quantized LLMs. Section 3.1 introduces the concept of CPU-augmented quantized LLM while Section 3.2 and Section 3.3 explore its opportunities and challenges.
Figure 2: CPU-augmented inference system for quantized LLMs.
@@ 3.1 Concept
Goal. We aim to utilize CPU memory as a means to enhance the quality of quantized LLMs without additional GPU memory costs. In model quantization, there is an inherent trade-off between size and quality: the more memory you allocate, the higher the quality you can achieve. When deploying quantized models in resource-constrained environments, it is common practice to explore this trade-off and determine the optimal bitwidth under the target GPU memory budget. This process can be a simple selection of a bitwidth to be uniformly applied to all parameters, or a more fine-grained allocation of bitwidths, such as layer-wise [15, 14, 8] or channel-wise [60, 57]. CPU memory may come into play once this process is complete, enabling further improvements without consuming additional GPU memory.
@@
Key Research Question. A key research question in designing an effective CPU-augmented inference system for quantized LLMs is determining a proper subset of residuals, or mask $\mathbf{M}$. A good mask $\mathbf{M}$ should: 1) select portions of the residuals that contribute most to improving model quality within the given bandwidth constraints, and 2) maintain a structured form that minimizes indexing overhead, thereby maximizing the effective transfer of residuals, while ensuring efficient processing on the GPU.
@@
Figure 4: Quantization error reduction trends observed when replacing the input channels of quantized weights with FP16 values sequentially in sorted order (solid lines) and random order (dotted lines) for the 16th decoder block of Llama-3-8B-Instruct. The distribution of activation magnitudes in sorted order is also shown (black lines).
Figure 5: (a) Distribution of activation outliers (top 5%) and (b) Recall rate of static analysis-based outlier identification for the true top 1% and 5% outliers, across 100 decoding steps. The 16th decoder block’s down projection of Llama-3-8B-Instruct model is used for profiling.
@@ 3.2 Opportunity: Not All Residuals Are Equally Important
Indeed, compensating for errors in salient channels using the corresponding residuals is highly effective. Figure 4 illustrates how the quantization error, defined as the mean square error between the computation result with FP16 weights (${\mathbf{W}}\mathbf{x}$) and quantized weights ($\widehat{\mathbf{W}}\mathbf{x}$), is reduced by sequentially replacing the input channels of the quantized weight with their corresponding FP16 values. For this analysis, 3-bit and 4-bit models of the Llama-3-8B-Instruct model [16], quantized using two state-of-the-art methods, AWQ [35] and SqueezeLLM [30], are evaluated. Specifically, the 7th decoding step is analyzed with the input prompt: "What are large language models?". All four linear layers in the 16th decoder block are included in this evaluation. The quantization errors for both 3-bit and 4-bit models are rapidly reduced when progressively compensating for channels in descending order of the corresponding activation magnitudes (solid red and solid blue lines). This trend closely follows the activation magnitude distribution (solid black line), which represents the sorted activation magnitudes in descending order. In contrast, the reduction in quantization error is significantly slower when input channels are compensated in random order, as shown by the dotted lines. This highlights the importance of prioritizing salient channels based on activation magnitude.
@@ 3.3 Challenge: Dynamic Nature of Activation Outliers
Identifying salient channels is challenging because the distribution of activation outliers changes dynamically by nature. While it is possible to infer salient channels by statically analyzing activation value statistics on a small calibration set [10, 32, 26, 35], such static approaches are suboptimal. Figure 5(a) shows the distribution of activation outliers—defined as activations with the top 5% magnitudes—in the down projection layer of the 16th decoder block of the Llama-3-8B-Instruct model over 100 decoding steps, again using the input prompt: "Large language models are". For better visibility, only the first 1024 channels are displayed. While some channels (e.g., Channel 421, highlighted by an arrow) consistently exhibit high activation magnitudes and remain persistent outliers, the distribution of activation outliers generally shows significant irregularity across different decoding steps.
@@
4 QDEC Design
Figure 6: QDEC overview.
@@ 4.1 Overview
In this section, building on the opportunity outlined in Section 3.2 and addressing the challenge described in Section 3.3, we propose QDEC, a CPU-augmented inference system for quantized LLMs that performs decoding with dynamic error compensation. Figure 6 presents an overview of QDEC. During the decode phase, QDEC augments each linear layer, essentially a GEMV operation, with dynamic error compensation. Specifically, to produce the final output ($\mathbf{o}$), QDEC adds an error compensation term ($\mathbf{o_{\text{dec}}}$) to the result of the base GEMV, $\mathbf{o_{\text{b}}}$, which is obtained by multiplying the input vector ($\mathbf{x}$) with the quantized weights ($\mathbf{\widehat{W}}$). $\mathbf{o_{dec}}$ is computed by multiplying the input vector with a subset of weight residuals selectively fetched from the CPU. This selection of weight residuals is performed dynamically, fully accounting for the changing distribution of input vectors. To maximize the number of residual values fetched under PCIe bandwidth constraints, QDEC stores and retrieves a quantized version of the residuals ($\mathbf{\widehat{R}}$), comprising the quantized values $Q_{r}(\mathbf{R})$ and associated metadata, instead of the full-precision residuals ($\mathbf{R}$). Here, $Q_{r}$ is a quantizer function for residuals that maps full-precision values to low-bit representations. Note that $Q_{r}$ is distinct from the quantizer function used for the original weight quantization, or base quantizer function, $Q_{b}$.
@@
The following sections provide details of each component of QDEC. Section 4.2 explains how QDEC performs the residual quantization. Section 4.3 details the GPU implementation of the dynamic error compensation. Section 4.4 illustrates how QDEC configures system parameters, including $k$, the number of channels to compensate.
@@ 4.2 Residual Quantization
Figure 7 depicts the quantization scheme for the residuals. QDEC employs 4-bit quantization for each output channel (i.e., column) of the residuals. To minimize metadata, symmetric uniform quantization is used. This approach requires only a single scalar scale factor as metadata for each output channel. The residual quantizer for the $i$-th output channel ($Q_{r,i}$) is defined as:
@@ 4.3 Efficient Implementation of Dynamic
The top priority in implementing dynamic error compensation is ensuring low latency, allowing its execution to remain hidden within the base GEMV execution time. To fetch a sufficient number of residual channels within this short time window, QDEC introduces three key software optimization strategies: 1) zero-copy residual fetching, 2) fast approximate Top-K for channel selection, and 3) kernel fusion.
@@
Zero-Copy Residual Fetch. QDEC leverages CUDA zero-copy [41], instead of commonly used API functions such as cudaMemcpy() or cudaMemcpyAsync(), to fetch the residuals from CPU. These APIs rely on the direct memory access (DMA) engine for data transmission, which is efficient at transferring large data blocks but suboptimal for smaller transfers due to the DMA setup overheads. Fetching residuals, however, falls into the latter category. The granularity of residual fetching occurs at the row level of the quantized residual matrix. With 4-bit quantization and typical row lengths of a few thousand to tens of thousands, each data block transfer is only a few tens of KBs. For optimal PCIe bandwidth utilization, the data block size should ideally be at least a few hundred KBs (e.g., 256 KB) [40, 46]. In zero-copy access, the GPU directly sends cacheline-sized memory requests, making it suitable for fine-grained data access. While zero-copy access has the disadvantage of occupying GPU cores to generate memory requests—potentially slowing down other concurrently running kernels—this is not a major issue for QDEC. The concurrent kernel for our case, the base GEMV, is typically memory-bound, so using fewer cores for this kernel is unlikely to have significant impact on its execution time.
Figure 8: Fast approximate Top-K operation of QDEC.
Figure 9: Bucket boundaries used in the approximate Top-K operation of QDEC.
@@
Fast Approximate Top-K for Channel Selection. For channel selection, QDEC employs an approximate Top-K method that is fast and GPU-friendly, while maintaining sufficient precision. This approach avoids the latency overhead of an exact Top-K operation. Figure 8 explains the approximation by providing an example where 128 elements need be selected from a 4096-dimensional input activation vector ($d_{in}=4096$, $k=128$). As shown in Figure 8(a), instead of performing one global selection over the entire input vector, QDEC conducts multiple local searches, one for each contiguous 1024-dimensional subvector, or chunk. For each chunk, $k_{chunk}$ elements are selected, where $k_{chunk}$ is $k$ divided by the number of chunks. In this example, we have four chunks, and $k_{chunk}$ is 32. The results from each local search are concatenated to form the final result. While this approach introduces approximation, it significantly reduces latency by eliminating the global synchronization, as each local search can be handled independently by a separate thread block.
@@
For each local search, QDEC uses a variant of the bucket-based Top-K algorithm [3]. Figure 8(b) illustrates this process. 1 First, the 1024 elements in a chunk are scattered into buckets based on their magnitudes, with bucket boundaries ($b^{k}_{0}$, $b^{k}_{1}$, …, $b^{k}_{30}$). The number of buckets is set to 32, matching the number of threads in a warp, allowing for efficient thread-level parallelism so that each thread processes a different bucket. 2 Next, elements are gathered, starting from bucket 0, until the total count reaches $k_{chunk}$. 3 If the number of elements in the current bucket exceeds the remaining spots for $k_{chunk}$, as in the case of bucket 9 in Figure 8(b), random selection is used to fill the remaining spots. This random selection adds an additional layer of approximation but significantly reduces latency by avoiding exact sorting.
@@
Determining proper bucket boundaries is crucial for minimizing the approximation error introduced by random selection ( 3 in Figure 8(b)). QDEC profiles the distribution of activation values using a small calibration set and aims to set boundary values that balance Top-K accuracy with the ability to handle a broader range of values. Placing finer-grained buckets around the expected $k$-th largest value generally improves accuracy, but it limits the system’s ability to handle out-of-distribution values. To address this, QDEC uses an offline analysis to determine two key boundaries, $b^{k}_{0}$ and $b^{k}_{15}$, from which the other boundaries are inferred (as shown in Figure 9). Assume the distribution of activation values from the calibration set is $\mathbf{X}\in R^{N\times d_{in}}$, where $N$ is the size of the set. The boundary $b^{k}_{15}$ is set to the maximum of the $k$-th largest value across all vectors in $abs(\mathbf{X})$. The interval between 0 and $b^{k}_{15}$ is uniformly divided into 16 buckets ($b^{k}_{15}$, $b^{k}_{16}$, …, $b^{k}_{30}$), focusing on the range where the $k$-th largest value is most likely to occur. To handle out-of-distribution cases which can cause significant degradation in selection precision, QDEC assigns an additional 16 buckets for values beyond $b^{k}_{15}$. Specifically, $b^{k}_{0}$ is set to the maximum value in the entire $abs(\mathbf{X})$, and the range between $b^{k}_{0}$ and $b^{k}_{15}$ is also uniformly divided to form the remaining 16 buckets.
@@
Figure 11: Parameter tuning process for QDEC, assuming a total of 56 SMs and a target slowdown rate of 10%.

### v2

# DecDEC: A Systems Approach to Advancing Low-Bit LLM Quantization

@@ Abstract
Quantization of Large Language Models (LLMs) has recently gained popularity, particularly for on-device settings with limited hardware resources. While efficient, quantization inevitably degrades model quality, especially in aggressive low-bit settings such as 3-bit and 4-bit precision. In this paper, we propose DecDEC, an inference scheme that improves the quality of low-bit LLMs while preserving the key benefits of quantization: GPU memory savings and latency reduction. DecDEC stores the residual matrix—the difference between full-precision and quantized weights—in CPU, and dynamically fetches the residuals for only a small portion of the weights. This portion corresponds to the salient channels, marked by activation outliers, with the fetched residuals helping to correct quantization errors in these channels. Salient channels are identified dynamically at each decoding step by analyzing the input activations—this enables adaptation to the dynamic nature of activation distribution, thus maximizing the effectiveness of error compensation. We demonstrate the effectiveness of DecDEC by augmenting state-of-the-art quantization methods. For example, DecDEC reduces the perplexity of a 3-bit Llama-3-8B-Instruct model from 10.15 to 9.12—outperforming its 3.5-bit counterpart—while adding less than 0.0003% to GPU memory usage and incurring only a 1.7% inference slowdown on NVIDIA RTX 4050 Mobile.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) based on the Transformer architecture [59] have shown great potential to reshape our daily lives [7, 56, 43, 39, 55]. However, their deployment costs pose a significant challenge, as large model sizes increase memory requirements and latency, limiting their use cases [60]. Quantization is a promising solution for reducing the LLM deployment costs [69, 27]. By lowering the model precision, quantization addresses both memory limitations and inference latency. The importance of quantization is pronounced for on-device deployments, where strict memory budgets often make model compression mandatory rather than optional. These scenarios typically require careful tuning of quantization levels to achieve an optimal balance between model size and quality [15, 14, 61, 8, 58].
While quantization may allow gigantic LLMs to fit into small-memory devices, it often leads to model quality degradation due to the inevitable loss of information. This is especially true for low-bit settings, such as 3-bit and 4-bit quantization, which are often used to accommodate the parameter sizes of LLMs [30, 35, 19]. This raises a key research question that this paper addresses: given a quantized LLM configured with the best possible effort under the memory budget, is there a way to recover the quality loss caused by quantization?
@@
A clue for this problem comes from the well-known fact that not all channels in the weight matrix are equally important for quantization. Some channels, referred to as salient channels, are more critical, primarily due to the presence of activation outliers [10, 6, 24]. When certain input activation values are particularly large, the quantization errors in the corresponding weight channels—those multiplied by these large activation values—are amplified, making the channels salient. By identifying these channels and selectively fetching error compensation terms from CPU memory, we can maximize the quality boost while minimizing data transfer.
Hence, the identification of the salient channels is essential for overcoming the bandwidth limitation, allowing for the transfer of only impactful information. Previous works that attempt to improve quantization algorithms by addressing activation outliers analyze the activation value distribution on a calibration data to predetermine the salient channels [10, 32, 26, 35]. This approach is suboptimal as it statically designates certain channels as salient throughout the inference, while the real distribution of activation values—and thus the distribution of salient channels—changes dynamically at each decoding step. Accounting for this dynamic nature is crucial for the accurate identification of salient channels.
Thus, in this paper, we introduce DecDEC (Decoding with Dynamic Error Compensation), an inference scheme for quantized LLMs that dynamically identifies salient channels and compensates for quantization errors in these channels, in real time. DecDEC enhances model quality while fully preserving the two main benefits of quantization: GPU memory savings and latency reduction. To achieve this, DecDEC stores the residuals of the quantized weight matrices in CPU memory, fetching only the parts that correspond to the dynamically identified salient channels for error compensation. Although small in size, these residuals provide a significant quality boost. This dynamic error compensation is performed concurrently with inference by an optimized GPU kernel, ensuring that all additional operations are seamlessly integrated into the existing workflow, minimizing inference slowdown. Below are the key contributions of our work:
@@
  • We present DecDEC, an inference scheme that enhances quantized LLMs by dynamically identifying salient channels and compensating quantization errors in them.
  • We introduce a tuner for DecDEC that recommends system parameters to satisfy a target latency bound.
  • We evaluate DecDEC across five different consumer-grade GPUs, demonstrating significant quality improvement with minimal memory and latency overhead.
@@ 2.2 LLM Quantization
Quantization for LLMs—a popular compression technique that reduces both memory usage and inference latency—can be categorized into two main types [69, 27]: weight-activation quantization [65, 10, 49, 64, 36, 33] and weight-only quantization [19, 35, 30, 12, 32, 9, 57, 17, 11, 29]. Each are suited to different inference scenarios. Weight-activation quantization is primarily used in datacenter settings, where both memory and computational costs must be minimized to improve throughput. Quantizing both weights and activations allows for the efficient use of low-precision arithmetic units (e.g., INT4, INT8, FP8) available on modern GPUs [1]. In contrast, for on-device inference, weight-only quantization is the preferred approach [69, 30]. In this apporach, quantized weights are loaded from memory and dequantized on-the-fly to full precision (i.e., FP16), before being multiplied with the full-precision activations [44]. Though it only reduces memory traffic, this is sufficient to speed up on-device inference where memory is the bottleneck, as discussed in Section 2.1.
Weight-only quantization can be further divided into two sub-categories: quantization-aware training (QAT) [36, 11, 29], and post-training quantization (PTQ) [19, 35, 30, 12, 32, 9, 57, 17]. While QAT yields better results by retraining to reduce quantization errors, its cost makes it impractical for many end-users [30, 69]. As a result, PTQ—requiring no retraining—has become the preferred method for on-device LLM inference. Therefore, in this paper, we specifically focus on weight-only PTQ for on-device inference.
@@ 3 Augmenting Quantized LLM with CPU Memory
In this section, we propose leveraging CPU memory—an often underutilized resource in LLM inference, as GPUs are the de facto standard processors for this workload—as a means to augment quantized LLMs. Section 3.1 introduces the concept of CPU-augmented quantized LLMs while Section 3.2 and Section 3.3 explore its opportunities and challenges.
Figure 2: CPU-augmented inference for quantized LLMs.
@@ 3.1 Concept
Goal. We aim to leverage CPU memory to improve quantized LLM quality without additional GPU memory costs. Quantization trades model quality for a smaller memory footprint. In constrained settings, practitioners optimize this trade-off within a fixed GPU memory budget by selecting a uniform bitwidth or applying fine-grained strategies such as layer-wise [15, 14, 8] or channel-wise [61, 58] allocation. Our goal is to then further improve quality post hoc by utilizing CPU memory, without increasing GPU memory usage.
@@
Key Research Question. A key research question in designing an effective CPU-augmented inference system for quantized LLMs is determining a subset of residuals, or mask $\mathbf{M}$. A good mask $\mathbf{M}$ should: 1) select portions of the residuals that contribute most to improving model quality within the bandwidth constraints, and 2) maintain a structured form that minimizes indexing overhead and ensures effective residual transfer, while enabling efficient processing on the GPU.
@@ 3.2 Opportunity: Not All Residuals Are Equally Important
Indeed, compensating for errors in salient channels using the corresponding residuals is highly effective. Figure 4 illustrates how the quantization error, defined as the mean squared error between the computation result with FP16 weights (${\mathbf{W}}\mathbf{x}$) and quantized weights ($\widehat{\mathbf{W}}\mathbf{x}$), is reduced by sequentially replacing the input channels of the quantized weight with their corresponding FP16 values. For this analysis, we evaluate 3-bit and 4-bit versions of the LLaMA-3-8B-Instruct model [16], quantized using the state-of-the-art method AWQ [35], with a text sample from the C4 dataset [47] as the input prompt. All four linear layers in the 8th, 16th, and 24th decoder blocks are included in this evaluation. The quantization errors for both the 3-bit and 4-bit models drop rapidly when we progressively compensate for channels in descending order of their activation magnitudes (solid red and solid blue lines). This trend closely follows the activation magnitude distribution (solid black line), which represents the sorted activation magnitudes in descending order. In contrast, the reduction in quantization error is significantly slower when input channels are compensated in random order, as shown by the dotted lines. This highlights the importance of prioritizing salient channels based on activation magnitude.
Figure 4: Quantization error reduction trends observed when replacing the input channels of quantized weights with FP16 values sequentially in sorted order (solid lines) and random order (dotted lines) for the 8th, 16th and 24th decoder block of Llama-3-8B-Instruct. The distribution of activation magnitudes in sorted order is also shown (black lines).
Figure 5: (a) Distribution of activation outliers (top 5%) and (b) Recall rate of static analysis-based outlier identification for the true top 1% and 5% outliers, across 100 decoding steps. The down projection layers in the 8th, 16th, and 24th decoder blocks of Llama-3-8B-Instruct model are used for profiling.
@@ 3.3 Challenge: Dynamic Nature of Activation Outliers
Identifying salient channels is challenging because the distribution of activation outliers changes dynamically by nature. While it is possible to infer salient channels by statically analyzing activation value statistics on a small calibration set [10, 32, 26, 35], such static approaches are suboptimal. Figure 5(a) shows the distribution of activation outliers—defined as activations with the top 5% magnitudes—in the down projection layer of the 8th, 16th, and 24th decoder blocks of the Llama-3-8B-Instruct model over 100 decoding steps, using a text sample for C4 dataset as the input prompt [47]. For visibility, only the first 512 channels are shown. While some channels (e.g., Channel 306 in 24th block, highlighted by an arrow) consistently exhibit high activation magnitudes and remain persistent outliers, the outlier distribution generally shows significant irregularity across decoding steps.
@@
4 DecDEC Design
Figure 6: DecDEC overview.
@@ 4.1 Overview
In this section, building on the opportunity outlined in Section 3.2 and addressing the challenge described in Section 3.3, we propose DecDEC, a CPU-augmented inference system for quantized LLMs that performs decoding with dynamic error compensation. Figure 6 presents an overview of DecDEC. During the decode phase, DecDEC augments each linear layer, essentially a GEMV operation, with dynamic error compensation. To produce the final output $\mathbf{o}$, DecDEC adds an error compensation term $\mathbf{o_{\text{dec}}}$ to the base GEMV result $\mathbf{o_{\text{b}}=\widehat{W}x}$. $\mathbf{o_{dec}}$ is computed by multiplying the input vector with a subset of weight residuals selectively fetched from the CPU. This selection is performed dynamically, fully accounting for the variability of input vectors. To maximize the number of residual values fetched under PCIe bandwidth constraints, DecDEC stores and retrieves a quantized version of the residuals ($\mathbf{\widehat{R}}$), comprising the quantized values $Q_{r}(\mathbf{R})$ and associated metadata, instead of the full-precision residuals ($\mathbf{R}$). Here, $Q_{r}$ is a quantizer that maps full precision residuals to low-bit form and differs from the base quantizer used for the weights, $Q_{b}$.
@@
The following sections provide details of each component of DecDEC. Section 4.2 explains how DecDEC performs the residual quantization. Section 4.3 details the GPU implementation of the dynamic error compensation. Section 4.4 illustrates how DecDEC configures system parameters, including $k$, the number of channels to compensate.
@@ 4.2 Residual Quantization
Figure 7 depicts the quantization scheme for the residuals. DecDEC employs 4-bit quantization for each output channel (i.e., column) of the residuals. To minimize metadata, symmetric uniform quantization is used. This approach requires only a single scalar scale factor as metadata for each output channel. The residual quantizer for the $i$-th output channel ($Q_{r,i}$) is defined as:
@@ 4.3 Efficient Implementation of Dynamic
The top priority in implementing dynamic error compensation is ensuring low latency, allowing its execution to remain hidden within the base GEMV execution time. To fetch a sufficient number of residual channels within this short time window, DecDEC introduces three key software optimization strategies: 1) zero-copy residual fetching, 2) fast approximate Top-K for channel selection, and 3) kernel fusion.
Figure 8: Fast approximate Top-K operation of DecDEC.
Figure 9: DecDEC ’s approximate Top-K bucket boundaries.
@@
Zero-Copy Residual Fetch. DecDEC leverages CUDA zero-copy [41], instead of commonly used API functions such as cudaMemcpy() or cudaMemcpyAsync(), to fetch the residuals from CPU. These APIs rely on the direct memory access (DMA) engine for data transmission, which is efficient at transferring large data blocks but suboptimal for smaller transfers due to the DMA setup overheads. Fetching residuals, however, falls into the latter category. The granularity of residual fetching occurs at the row level of the quantized residual matrix. With 4-bit quantization and typical row lengths of a few thousand to tens of thousands, each data block transfer is only a few tens of KBs. For optimal PCIe bandwidth utilization, the data block size should ideally be at least a few hundred KBs (e.g., 256 KB) [40, 46]. In zero-copy access, the GPU directly sends cacheline-sized memory requests, making it suitable for fine-grained data access. While zero-copy access has the disadvantage of occupying GPU cores to generate memory requests—potentially slowing down other concurrently running kernels—this is not a major issue for DecDEC. The concurrent kernel for our case, the base GEMV, is typically memory-bound, so using fewer cores for this kernel is unlikely to have significant impact on its execution time.
@@
Fast Approximate Top-K for Channel Selection. For channel selection, instead of an exact Top-K operation, DecDEC employs an approximate Top-K method that is fast and GPU-friendly while maintaining precision. Figure 8 illustrates the approach using an example where 128 elements are selected from a 4096-dimensional activation vector ($d_{in}=4096$, $k=128$). As shown in Figure 8(a), instead of a single global selection, DecDEC partitions the input into four contiguous 1024-dimensional chunks and performs a local Top-$k_{chunk}$ selection within each chunk. In this example, $k_{chunk}=k/4=32$. The locally selected elements are then concatenated to form the final result. Although this introduces approximation, it significantly reduces latency by avoiding global synchronization—each local selection is handled independently by a thread block. A larger chunk size can improve efficiency but may increase the approximation error; we set the chunk size to 1024 to balance this trade-off effectively.
@@
For each local selection, DecDEC uses a variant of the bucket-based Top-K algorithm [2]. Figure 8(b) illustrates this process. 1 First, the 1024 elements in a chunk are scattered into buckets based on their magnitudes, with bucket boundaries ($b^{k}_{0}$, $b^{k}_{1}$, …, $b^{k}_{30}$). The number of buckets is set to 32, matching the number of threads in a warp, allowing for efficient thread-level parallelism so that each thread processes a different bucket. 2 Next, elements are gathered, starting from bucket 0, until the total count reaches $k_{chunk}$. 3 If the number of elements in the current bucket exceeds the remaining spots for $k_{chunk}$, as in the case of bucket 9 in Figure 8(b), random selection is used to fill the remaining spots. This random selection adds an additional layer of approximation but significantly reduces latency by avoiding exact sorting.
@@
Determining proper bucket boundaries is crucial to minimize the approximation error introduced by random selection ( 3 in Figure 8(b)). DecDEC profiles the distribution of activation values using a small calibration set and aims to set boundary values that balance Top-K accuracy with the ability to handle a broader range of values. Placing finer-grained buckets around the expected $k$-th largest value generally improves accuracy, but limits the system’s ability to handle out-of-distribution values. To address this, DecDEC uses an offline analysis to determine two key boundaries, $b^{k}_{0}$ and $b^{k}_{15}$, from which the other boundaries are inferred (as shown in Figure 9). Let the distribution of activation values from the calibration set be $\mathbf{X}\in R^{N\times d_{in}}$, where $N$ is the size of the set. The boundary $b^{k}_{15}$ is set to the maximum of the $k$-th largest value across all vectors in $abs(\mathbf{X})$. The interval between 0 and $b^{k}_{15}$ is uniformly divided into 16 buckets ($b^{k}_{15}$, $b^{k}_{16}$, …, $b^{k}_{30}$), focusing on the range where the $k$-th largest value is most likely to occur. To handle out-of-distribution cases which can cause significant degradation in selection precision, DecDEC assigns an additional 16 buckets for values beyond $b^{k}_{15}$. Specifically, $b^{k}_{0}$ is set to the maximum value in all $abs(\mathbf{X})$, and the range between $b^{k}_{0}$ and $b^{k}_{15}$ is also uniformly divided to form the remaining 16 buckets.
@@
Figure 11: Parameter tuning process for DecDEC, assuming a total of 56 SMs and a target slowdown rate of 10%.